OpenStudy (anonymous):

find x: x+.75x=210 Please tell steps

OpenStudy (anonymous):

Combine like terms: 1x+.75x=210 1.75x=210 Divide each side by 1.75 1.75x/1.75=210/1.75 x=210/1.75 x=120

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you're welcome